64 identical drops each of capacity 5 micro facade combine to form a big drop. What is the capacity of big drop.

Grade:12

1 Answers

Mangesh Y S
7 years ago
Capacitor of big drop is 20micro farad.
This can be obtained by the expression
Cbig=n1/3 Csmall
=641/3 5
=4×5
=20micro farad